# Moral Alignment for LLM Agents

### v1

@@ 1 Introduction
We explore the proposed framework using an Iterated Prisoner’s Dilemma environment, in which we evaluate the effectiveness of fine-tuning based on intrinsic rewards as a mechanism for learning moral strategies as well as “unlearning” a selfish strategy. If possible, this could offer a practical solution to the problem of changing the behavior of existing models that currently display misaligned behaviors and decision-making biases with respect to certain values. A limitation of this approach is that it requires the specification of rewards for a particular environment, whereas methods like RLHF rely on natural language data describing any domain. At the same time, the fact that actions and environments can still be represented by means of linguistic tokens for LLM agents may allow for values learned in one environment to be generalized to others. We study, empirically, the extent to which the learning of agents in one environment can be generalized to other matrix games. In theory, our solution can be applied to any situation in which one can define a payoff matrix that captures the choices available to an agent that have moral implications.
@@ 2.3 Social Dilemma Games
In a single iteration of the game, the payoffs motivate each player to Defect due to the risk of facing an uncooperative opponent (i.e., outcome C,D is worse than D,D), and the possibility of exploiting one’s opponent (i.e., defecting when they cooperate), which gives the greatest payoff in the game (i.e., D,C is preferred over C,C). Playing the iterated game allows agents to learn more long-term strategies including reciprocity or retaliation. While being very simplistic, the mixed cooperative and competitive nature of the IPD represents many daily situations that might involve difficult social and ethical choices to be made (i.e., moral dilemmas). This is why it has been extensively used for studying social dilemmas in traditional RL-based agents (Bruns, 2015; Hughes et al., 2018; Anastassacos et al., 2020; McKee et al., 2020; Leibo et al., 2021) and, more recently, utilized as a training environment for moral alignment of agents in particular (Tennant et al., 2023a; 2024).
@@ 3.1 Agent and Environment
The LLM agent and an opponent play a repeated Iterated Prisoner’s Dilemma game. At each time step, the model receives a prompt containing a description of the IPD game, including a state that contains the history of the opponent’s single previous move. Within the MDP framework, each player’s current action affects the game’s state at the next time step.
@@ 4.3 Learning and Unlearning the Selfish Strategy on the IPD
Figure 3: Action types played by the LLM agent during different types of fine-tuning on the Iterated Prisoner’s Dilemma (IPD) game a) vs a TFT agent, and b) vs an LLM agent (i.e., two LLMs being fine-tuned at once). For each episode, we plot the actions of the LLM player $M$ given the last move of their opponent $O$. In these “unlearning” experiments, the reward function changes from the IPD Game payoffs to a moral intrinsic reward (Deontological or Utilitarian) at episode 500.
@@ 5 Generalization to Moral Choices in Other Environments
After fine-tuning the models with moral reward, we evaluate each one through 10 episodes: each episode begins with a randomly generated state and proceeds through 5 interaction steps. We average the results across the 5 runs of each fine-tuned model. In this section, we present evaluations of models which were fine-tuned versus a static opponent. The results for models trained against another LLM show similar patterns - these are reported in the Appendix (Section 8.7).

### v3

@@ 1 Introduction
We explore the proposed framework using an Iterated Prisoner’s Dilemma environment, in which we evaluate the effectiveness of fine-tuning based on intrinsic rewards as a mechanism for learning moral strategies as well as “unlearning” [^2] We note that by “unlearning” we refer to re-prioritizing certain principles in an agent’s decision-making. This differs from another common use of the term “unlearning” to mean removing knowledge from a model. a selfish strategy. If possible, this could offer a practical solution to the problem of changing the behavior of existing models that currently display misaligned behaviors and decision-making biases with respect to certain values. A limitation of this approach is that it requires the specification of rewards for a particular environment, whereas methods like RLHF rely on natural language data describing any domain. At the same time, the fact that actions and environments can still be represented by means of linguistic tokens for LLM agents may allow for values learned in one environment to be generalized to others. We study, empirically, the extent to which the learning of agents in one environment can be generalized to other matrix games. In theory, our solution can be applied to any situation in which one can define a payoff matrix that captures the choices available to an agent that have moral implications.
@@ 2.3 Social Dilemma Games
In a single iteration of the game, the payoffs motivate each player to Defect due to the risk of facing an uncooperative opponent (i.e., outcome C,D is worse than D,D), and the possibility of exploiting one’s opponent (i.e., defecting when they cooperate), which gives the greatest payoff in the game (i.e., D,C is preferred over C,C). Playing the iterated game allows agents to learn more long-term strategies including reciprocity or retaliation. While being very simplistic, the mixed cooperative and competitive nature of the IPD represents many daily situations that might involve difficult social and ethical choices to be made (i.e., moral dilemmas). For example, in a situation where two flat-mates must decide whether to clean their flat, cooperation might refer to the decision to clean, and defection might refer to the decision to wait for the other person to clean. This is why it has been extensively used for studying social dilemmas in traditional RL-based agents (Bruns, 2015; Hughes et al., 2018; Anastassacos et al., 2020; McKee et al., 2020; Leibo et al., 2021) and, more recently, utilized as a training environment for moral alignment of agents in particular (Tennant et al., 2023a; 2024).
@@ 3.1 Agent and Environment
The LLM agent and an opponent play a repeated Iterated Prisoner’s Dilemma game. At each time step, the model receives a prompt containing a description of the IPD game, including a state that contains the history of the opponent’s single previous move (see Figure 6 in the Appendix). Within the MDP framework, each player’s current action affects the game’s state at the next time step.
@@ 4.3 Learning and Unlearning the Selfish Strategy on the IPD
Figure 3: “Unlearning” experiments, where the reward function changes from the IPD Game payoffs to a moral intrinsic reward (Deontological or Utilitarian) at episode 500. Action types played by the LLM agent during different types of fine-tuning on the Iterated Prisoner’s Dilemma (IPD) game a) vs a TFT agent, and b) vs an LLM agent (i.e., two LLMs being fine-tuned at once). For each episode, we plot the actions of the LLM player $M$ given the last move of their opponent $O$.
@@ 5 Generalization to Moral Choices in Other Environments
After fine-tuning the models with moral reward, we evaluate each one through 10 episodes: each episode begins with a randomly generated state and proceeds through 5 interaction steps. We average the results across the 5 runs of each fine-tuned model. In this section, we present evaluations of models which were fine-tuned versus a static (i.e., TFT) opponent. The results for models trained against another LLM show similar patterns - these are reported in the Appendix (Section 8.7).